*** Linear Regression Analysis ***

Dependent Variable:   Weight Loss (in Pounds)

Independent variable: Exercise Time (in Minutes)

                                                                        Analysis of Variance

                                                                     Sum of            Mean             F           p

                                    Source       df           Squares          Square       Ratio    Value

-------------------------------------------------------------------------

Regression    1        ___(b)___       85.456    __ (e)__     .001

Residual __(a)__      25.678          __(d)__

                                    -------------------------------------------------------------------------

Total           11        ___(c)___

3%

  1. Degree of Freedom for Residual = ____

TYPE YOUR ANSWER HERE: ____

3%

  1. Sum of Squares Due to Regression = ____

TYPE YOUR ANSWER HERE: ____

3%

  1. Sum of Squares for Total = _____

TYPE YOUR ANSWER HERE: ____

3%

  1. Mean Square for Residual = _____

TYPE YOUR ANSWER HERE: ____

3%

  1. F Ratio = _____

TYPE YOUR ANSWER HERE: ____

4%

  1. Interpret the result of F test for this study

TYPE YOUR ANSWER HERE: ____
Answer #1

a.

Degree of Freedom for Residual = DF Total - DF Regression = 11 - 1 = 10

b.

Sum of Squares Due to Regression = Mean Square for Regression * DF Regression

= 85.456 * 1 = 85.456

c.

Sum of Squares for Total = Sum of Squares Due to Regression + Sum of Squares Due to Residual

= 85.456 + 25.678 =  111.134

d.

Mean Square for Residual = Sum of Squares Due to Residual / DF Residual

= 25.678 / 10

= 2.5678

e.

F Ratio = Mean Square for Regression / Mean Square for Residual

= 85.456 / 2.5678

= 33.27985

f.

P-value = 0.001

Since, p-value is less than 0.05 significance level, we reject null hypothesis H0 and conclude that there is strong evidence that true mean weight loss is different for at least one of exercise time.